find the roots using newton raphson. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

e^x+x=4
1
Expert's answer
2020-10-25T18:51:42-0400
"f(x)=e^x+x-4"

"f'(x)=e^x+1"

"x_{n+1}=x_n-\dfrac{f(x_n)}{f'(x_n)}"

Initial solution "x_0 =1"

"\begin{matrix}
 n & x_n & f(x_n) & x_{n+1} \\
 0 & 1 & -0.281718 & 1.075766\\
 1 & 1.075766 & 0.008003 & 1.073729 \\
 2 & 1.073729 & 0.000006 & 1.073729 \\
 
\end{matrix}"


"\varepsilon =\big|\dfrac{x_{n+1}-x_n}{x_n}\big|\cdot 100\%"

"\varepsilon =\big|\dfrac{1.075766-1}{1}\big|\cdot 100\%=7.5766\%"

"\varepsilon =\big|\dfrac{1.073729-1.075766}{1.075766}\big|\cdot 100\%=0.1894\%"

"\varepsilon =\big|\dfrac{1.073729-1.073729}{1.073729}\big|\cdot 100\%=0.000000\%"

"x=1.073729"


The root is "1.073729"